OpenStudy (anonymous):

Medal!! help! You conduct an experiment where you roll a die and record the results. You roll the die 1230 times and the results are shown in the table below: What is the experimental probability of rolling a "1" on your next roll? A. 0.1667 B. 0.1992 C. 0.245 D. 0.3275

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry its not C, i know that forsure. What do you guys think?

OpenStudy (anonymous):

@wolf1728 ?

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@blurbendy can you help me please?

OpenStudy (blurbendy):

wait, you're sure it's not C ?

OpenStudy (anonymous):

yep, i got that one wrong last time, so im checking through again to make sure i get the correct answer

OpenStudy (blurbendy):

and the frequency on the last one isn't 375, it's 0.375?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

its 256 my bad.

OpenStudy (anonymous):

so the last one on frequency is 256 lol.

OpenStudy (blurbendy):

ah! okay, then it's B experimental probability = # times event occurs/total number of events so you roll a 1, 245 times, and the total number of events is 245 + 172 + 219 + 201 + 137 + 256 = 1230 = 245/1230 = 0.19918699187 which is basically B

OpenStudy (anonymous):

thanksyou sorry for he mistake! even i was like what the ahahaha

OpenStudy (blurbendy):

haha, it's all good. anytime :)